Calculate the area bounded by y = x³, x = 0 and y = 4.

1
Expert's answer
2022-02-25T04:20:57-0500

Tracing the graph of "y=x^3" shows us that the area is bounded by "x=0" on the left, "y=4" from above and "y=x^3" from below, with two latter encountering at "x=4^{1\/3}". Therefore, the area is given by :

"S=\int_0^{\sqrt[3]4}(4-x^3)dx" this integral can be easily calculated :

"S=4\cdot\sqrt[3]{4} - \frac{1}{4}x^4|^{\sqrt[3]{4}}_{0}"

"S=4 \sqrt[3]{4}-\sqrt[3]{4}=3\cdot \sqrt[3]{4}"